Express as a trinomial
(U+3)^2
(U-8)^2
(2u+5)^2
(1-4u)^2
(U+2v)^2
(7u-3v)^2
(I've+6)^2
PleAse help Im terrible at algerbra I don't even know how to go about solving this

Answer by rothauserc(4718)

A trinomial is a polynomial with three terms, a polynomial can have variables, constants and exponents(only 0, 1, 2, 3, 4,.... are allowed)
:
(U+3)^2 = (U+3) * (U+3) = U^2 +6U +9
:
(U-8)^2 = (U-8) * (U-8) = U^2 -16U +64
:
(2u+5)^2 = (2u+5) * (2u+5) = 4u^2 +20u +25
:
(1-4u)^2 = (1-4u) * (1-4u) = 1 -8u +16u^2
:
(U+2v)^2 = (U+2v) * (U+2v) = U^2 +4Uv +4v^2
:
(7u-3v)^2 = (7u-3v) * (7u-3v) = 49u^2 -42uv +9v^2
:
(I've+6)^2 = (I've+6) * (I've+6) = I've^2 +12I've +36
: